A rectangle has a length of 15 cm and a width of 8 cm. Use the expression 2w + 2l to find the perimeter of the rectangle.

Answer by CharlesG2(834)

A rectangle has a length of 15 cm and a width of 8 cm. Use the expression 2w + 2l to find the perimeter of the rectangle.

length l is known and is 15 cm, width w is known and is 8 cm
P perimeter rectangle = 2w + 2l, just need to plug in the values

P = 2(15) + 2(8) = 30 + 16 = 46 cm